Please help.
Solve the system by addition.
5x-3y=13
4x-3y=11
Thanks
JMS

Answer by bucky(2189)

To solve by addition, one of the equations must have a term that is equal to the same term in
the other equation, but these terms have opposite signs.
.
You are given the two equations:
.
+5x-3y=+13
+4x-3y=+11
.
Notice that the term 3y appears in both equations, but they have the same sign. How about
if we multiply the bottom equation (all terms on both sides) by -1. When we do that the
pair of equations becomes:
.
+5x-3y=+13
-4x+3y=-11
.
Now you can add the two equations in vertical columns. When you do the -3y and the +3y
cancel each other out and the addition results in:
.
x = 2
.
Substitute that value of x into either of the original problems and solve for y. Let's make
that substitution into the first equation. Start with:
.
+5x - 3y = 13
.
Substitute -2 for x to get:
.
5(2) - 3y = 13
.
Multiply out the first term:
.
10 - 3y = 13
.
subtract 10 from both sides of this equation:
.
-3y = +3
.
Divide both sides by -3 to solve for y and you get:
.
y = 3/-3 = -1
.
In summary the answer to this problem is x = 2 and y = -1.
.
Hope this helps you to understand the problem and how to solve it.
.